Kimberly Young is saving to buy a house in five years. She plans to put 20 percent down at that time, and she believes that she will need $39,000 for the down payment. If Kimberly can invest in a fund that pays 8.00 percent annual interest, compounded quarterly, how much will she have to invest today to have enough money for the down payment?
Answer #1

We use the formula:
A=P(1+r/4)^4n
where
A=future value
P=present value
r=rate of interest
n=time period.

39,000=P*(1+0.08/4)^(4*5)

P=39,000/(1+0.08/4)^(4*5)

=$39,000*0.672971333

=$26245.88(Approx).